OpenStudy (anonymous):

why is this true? \[\frac{2}{8i}=-\frac i 4\]

OpenStudy (anonymous):

first off you can reduce, just like with real numbers then multiply top and bottom by \(i\)

OpenStudy (anonymous):

because i*i=-1

OpenStudy (anonymous):

or 1/i=-i

OpenStudy (anonymous):

Is our goal to get the imaginary number out of the denominator?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{2}{8i}=\frac{1}{4i}=\frac{1}{4i}\times \frac{i}{i}=\frac{i}{4i^2}=-\frac{i}{4}\]

OpenStudy (anonymous):

What is wrong if I say 1/i=i

OpenStudy (anonymous):

it is \(\frac{1}{i}=-i\)